# A Spatiotemporal, Quasi-experimental Causal Inference Approach to Characterize the Effects of Global Plastic Waste Export and Burning on Air Quality Using Remotely Sensed Data

### v2

@@ 4.1 Statistical Analysis Results
Figure 3 shows the estimated ADT curve and 95% bootstrap confidence interval (CI) from the main analysis. Positive ADT estimates provide evidence of post-China ban increases in PM2.5 concentrations over GPW sites, compared to those expected under business-as-usual. Thus, our results indicate that at open dump sites in Indonesia, the monthly PM2.5 in 2018-2019 was higher than expected under business-as-usual (except for dump sites nowhere near port activity), with the effect ranging from -0.17–1.72 $\mu g/m^{3}$ (95% CI = (0.84, 2.66) for the maximum value) depending on the port proximity. As expected, we see smaller effects at locations with lower port proximity – the effect is not statistically significantly different from zero until $\delta=0.2$ (at which $\Psi(\delta)\geq 0.76$), above which it remains significant. The decrease of the effect curve between medium and high levels of port proximity ($\delta\in$ [0.7,1]) could reflect smaller increases in dumping/burning in densely populated areas very close to ports, where there is likely more government oversight and less open space. Integrating across all values of the estimated effect curve yields an average increase at dump sites of 1.14 $\mu g/m^{3}$ attributable to China’s ban on plastic imports, corresponding to a 4.4% increase in overall PM2.5 in 2018-2019 compared to 2012-2017. Note that if we only consider dump sites above the 20th quantile of port proximity (where the effect is statistically significant), the average increase is 1.38 $\mu g/m^{3}$, corresponding to a 5.1% increase at those sites.
@@
Figure 3: Average dose-effect on the treated (ADT) estimates and bootstrap 95% confidence intervals for our case study, estimated using both our proposed multiply robust method and a conventional outcome regression. ADT estimates (y-axis) quantify the average change in PM2.5 ($\mu g/m^{3}$) post-China ban, compared to concentrations expected under business-as-usual, at open dump sites in Indonesia for a given port proximity (x-axis). The pointwise confidence intervals are generated using our spatial weighted bootstrap with Matérn correlation range=26.8km (as observed in the empirical residuals).
@@
The results from our proposed multiply-robust method are similar to those produced using an outcome regression alone (Figure 3). But variations in both the curve (point estimate) and the confidence interval indicate that it is still important to guard against misspecification of the outcome model by incorporating propensity scores. By contrast, the confounder-naive results look very different (Figure S6): the mean effect ranges from -6.4 to 7.2 $\mu g/m^{3}$. This divergence from our main result is unsurprising, but demonstrative.
The results of using different values as the range parameter for the bootstrap weights’ spatial correlation are shown in Figure S7. A range of 10km produces confidence intervals 74% the width, on average, of those in the main analysis (with a range of 26.8km). Ignoring spatial dependence entirely in the weighted bootstrap narrows the confidence intervals by more than a factor of two (on average, they are 45% the width of those in the main analysis). This demonstrates that accounting for residual spatial correlation is essential to avoid under-estimation of uncertainty in this application.
@@
Figure 4: Pre-trends test using 2017 as the “treated” period; sampling bootstrap weights with spatial correlation (range=26.8km).
@@
As shown in Figure 4, we did not observe a significant increase in PM2.5 when using 2017 as the “treated” period (to check for pre-trends / policy anticipation) – in fact, the mean effect is negative, indicating lower PM2.5 in 2017 compared to preceding years. Using 2016 & 2017 as the “treated” period (Figure S8) yields the same conclusion.
Including GPW site area as a covariate in our sensitivity analysis did not substantially change the results. The multiply-robust curve in Figure S9 is slightly less peaked (its maximum value is 1.45 rather than 1.72), but the confidence interval widths are practically identical.
4.2 Post-Hoc Exploratory Data Analysis Results
@@ 5 Discussion
In this case study, we observed a 4.4% increase in PM2.5 at dump sites in Indonesia in 2018-2019 compared to 2012-2017. This effect is on a similar scale to the results of Unfried and Wang (2022), who assessed the domestic impacts of China’s plastic waste import ban and estimated an average reduction of 3.7 $\mu g/m^{3}$, corresponding to 8.5% reduction of overall PM2.5 concentrations in China as a result of their 2018 ban. It makes sense that our observed effect would be smaller, given that Indonesia was only one of the countries to which waste that would otherwise have gone to China was diverted. Unfried and Wang (2022) similarly did not observe an anticipation effect in mid-2017, when China announced its upcoming ban to the World Trade Organization.
@@
Second, while our interest is primarily on air pollution generated by the burning of waste, it is possible that some of the observed increase in PM2.5 at open dump sites is attributable to the transportation of this waste, for instance from diesel trucks making more trips to dump sites and, for sites close to the coast, from cargo ships. Our post-hoc analysis illustrating the uptick in active fires overlapping with dump sites offers some evidence that at least part of the observed effect could plausibly be due to increases in waste burning. Unfortunately, speciated PM2.5 estimates, which might enable the disentangling of emissions from waste burning and transportation, tend to be available only at coarse spatiotemporal resolutions globally and rely more heavily on models (European Commission, Joint Research Centre, 2023), which is likely to frustrate the application of quasi-experimental statistical methods. For smaller scale studies, Islam et al. (2022) demonstrated the viability of measuring a molecular tracer of plastic burning with in-situ monitors. Alternatively, Scott et al. (2023) applied machine learning-based image classification to identify smoke plumes in high-resolution (3m x 3m) satellite imagery over the largest municipal landfill in the Maldives. However, in addition to being computationally intensive, this approach is not directly translatable into air pollution concentrations, which are easier to link to the rich literature on health impacts, and hence tend to be more relevant for policy makers.
@@
A more general limitation of our approach is that we cannot estimate the effects of distributed waste burning, i.e., at the household level or at communal dump sites that were too small to be detected by GPW. Future research, and likely investment in ground-level air quality monitors (and/or lower cost sensors), will be needed to estimate these distributed burning emissions, which may ultimately pose greater risks to public health due to higher frequency of burning, greater proximity to people, and lower dispersive dilution of air pollution at the ground level (Pathak et al., 2024) compared to taller plumes of smoke resulting from larger fires at larger dump sites.
Despite its limitations, this study provides strong evidence that the export of plastic waste to Indonesia, and subsequent open burning of import-related waste in Indonesia, has resulted in higher concentrations of PM2.5 – at least near large open dump sites. This motivates future policies to (a) reduce the amount of plastic waste being exported from the Global North to the Global South and (b) reduce the amount of waste being openly burned. Both will require substantial investments in waste management infrastructure and/or dramatic shifts in the amount of plastic waste being generated. With respect to waste burning, it has been observed that simply prohibiting open burning is seldom effective (Pathak et al., 2024), though burning at major sites may be easier to monitor and restrict (Scott et al., 2023).
On the global stage, negotiations by the United Nations are currently underway to develop a Global Plastic Treaty – with an international legally binding instrument on plastic pollution (United Nations Environment Programme (2024), UNEP) – under the Basel Convention on the Control of Transboundary Movements of Hazardous Wastes and Their Disposal. The effectiveness of this and future policies, at least for improving air quality but potentially also for other kinds of environmental outcomes, can in turn be evaluated using methods similar to those presented in this paper.
Acknowledgements
This work was supported by an NSF Graduate Research Fellowship (EMC), NIH award K01ES032458 (RCN), and the Harvard Climate Change Solutions Fund (RCN). The computation for this paper was performed on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing at Harvard University.
@@ Appendix A Estimation Details
A.1 Nuisance Function Hyperparameters
For each of the nuisance functions ($\mu_{1},\mu_{0},\pi_{B},\pi_{D}$), we use the Ranger algorithm (from the R caret package) with num.trees=500. For the regressions (all but $\pi_{B}$, which is a binary classification), we use min.node.size=5 and select the minimum value of mtry (the number of variables to possibly split at in each node) beyond which the $R^{2}$ and RMSE do not substantially improve (a difference within 0.01 for $R^{2}$ and 0.1 for RMSE). For $\pi_{B}$, we limit mtry such that the empirical probabilities do not get too large ($>$0.99) or too small ($<$0.001).
To prevent overfitting, we use cross-validation with k=2 folds to select mtry. For $\mu_{0}$, we sample observations into the folds by year block (2012 & 2013, 2014 & 2015, 2016 & 2017) to emulate this model’s extrapolation to the $B=1$ years. For $\pi_{D}$, we sample observations by GPW site to account for $D$ being a static variable in our dataset, even though $\mathbf{X}$ contains some time-varying covariates. For $\mu_{1}$ and $\pi_{B}$, we randomly sample observations from $\mathcal{O}_{B=1}$ and $\mathcal{O}$ respectively.
A.2 Generalized Propensity Score (GPS) Estimation
@@
[…]
@@ Appendix B Supplementary Figures and Tables
Figure S5: Main analysis run with LLKR bandwidth selected on all data (with repeated observations at each dump site).
Figure S6: The confounder-naive results. Sampling bootstrap weights with spatial correlation (range=26.8km).
Figure S7: Comparing the results when the bootstrap weights have spatial autocorrelation range 0, 20, and 26.8 km respectively. The top righthand plot shows the Matérn correlation decay versus distance, using these different values for the range parameter.
Figure S8: Pre-trends test for 2016-2017. Sampling bootstrap weights with spatial correlation (range=26.8km).
Figure S9: Sensitivity analysis including GPW site area as a covariate. Sampling bootstrap weights with spatial correlation (range=26.8km).
@@ Appendix C Simulations Exploring the Spatial Weighted Bootstrap
In this appendix, our primary aim is to evaluate the characteristics of the confidence intervals generated by our Spatial Weighted Bootstrap procedure, relative to those generated by a Non-spatial Weighted Bootstrap. Here, calculation of the ADT curve’s bias and Mean Absolute Error (MAE) is an intermediate step to better understand the confidence intervals, as Hettinger et al. (2025) present extensive investigation of a very similar ADT estimator.
@@ C.1 Simulating the Outcome
$\tilde{Y}_{it}=\mu_{it}+\sigma_{\text{sp}}\epsilon_{\text{sp}}+\sigma_{\text{ind}}\epsilon_{\text{ind}}$
where $\epsilon_{\text{sp}}\sim MVN(0,\Sigma)$, where the correlation function to generate the entries of $\Sigma$ is $Corr(m;L)=exp(-m/L)$ [Exponential form] and $\epsilon_{\text{ind}}\sim N(0,1)$. The parameters $L$ (the true spatial range), $\sigma_{\text{sp}}$, and $\sigma_{\text{ind}}$ are constants, which we specify to obtain four distinct simulation scenarios:
@@ C.2 Estimating the Effect Curve and Confidence Interval
When implementing our ADT estimation procedure on each simulated dataset, we correctly specify the outcome models’ mean function but estimate the propensity score models $\pi_{B}$ and $\pi_{D}$ using Ranger (the same algorithm as in our main analysis).
When applying the Spatial Weighted Bootstrap, we fit both Exponential and Power covariance functions (the latter has correlation function $Corr(m;a,z)=1-zm^{a}$) to the empirical variogram of the residuals from the LLKR stage and use whichever model has the smaller RMSE to generate the spatial weights. As noted in Section 3.4.2, the spatial structure of the LLKR residuals is different than the spatial structure of the outcome (here, $\tilde{Y}$), so traditional notions of correctly vs incorrectly specifying the spatial correlation model do not directly apply here.

### v3

@@ 4.1 Statistical Analysis Results
Figure 4: Average dose-effect on the treated (ADT) estimates and bootstrap 95% confidence intervals for our case study, estimated using both our proposed multiply robust method and a conventional outcome regression. ADT estimates (y-axis) quantify the average change in PM2.5 ($\mu g/m^{3}$) post-China ban, compared to concentrations expected under business-as-usual, at open dump sites in Indonesia for a given port proximity (x-axis). Note that our port proximity metric is in quantile form, so the observed support of the exposure along the x-axis of these plots is constant. The pointwise confidence intervals are generated using our Spatial Weighted Bootstrap.
Figure 4 shows the estimated ADT curve and 95% bootstrap confidence interval (CI) from the main analysis. Positive ADT estimates provide evidence of post-China ban increases in PM2.5 concentrations at GPW sites, compared to those expected under business-as-usual. Thus, our results indicate that at open dump sites in Indonesia, the monthly PM2.5 in 2018-2019 was generally higher than expected under business-as-usual (except for dump sites nowhere near port activity), with the effect getting as large as 1.68 $\mu g/m^{3}$ (95% CI = [0.72, 2.48]) depending on the port proximity. As expected, we see smaller effects at locations with lower port proximity – the pointwise confidence intervals include zero until $\delta=0.38$. Interestingly, the effect curve peaks in the middle and decreases for high values of port proximity, including zero again at $\delta=0.81$ – our intuition is that this could reflect smaller increases in dumping/burning in densely populated areas very close to ports, where there is likely more government oversight and less open space. Integrating across all values of the estimated effect curve yields an average increase at dump sites of 0.86 $\mu g/m^{3}$ attributable to China’s ban on plastic imports, corresponding to a 3.4% increase in overall PM2.5 in 2018-2019 compared to 2012-2017. If we only consider dump sites where the confidence intervals do not include zero, the average increase is 1.51 $\mu g/m^{3}$, corresponding to a 5.3% increase at those sites.
@@
While similarity between the effect curve from our proposed multiply-robust method and that produced using an outcome regression alone (Figure 4) suggests a fairly strong outcome model, variations in the tails highlight the importance of guarding against misspecification of the outcome model by incorporating propensity scores. We also see that the generalized propensity score estimation is more uncertain, resulting in substantially wider confidence intervals for the multiply-robust method than for the outcome regression alone. For a more extreme comparison, the confounder-naive results are shown in Figure S6: the mean effect ranges from -6.4 to 8.5 $\mu g/m^{3}$. This divergence from our main result is unsurprising, but demonstrative.
The results of ignoring residual spatial correlation in the bootstrap stage are shown in Figure S7. The confidence intervals are, on average, 73% the width of the intervals obtained from Spatial Weighted Bootstrap. Note that this is for our multiply robust method, which also incorporates spatial-folds (all observations from a site sampled together) cross-fitting when estimating the generalized propensity score. For the outcome regression alone, the confidence intervals from the Non-Spatial Weighted Bootstrap are 52% the width of the intervals from the Spatial Weighted Bootstrap, on average. This demonstrates the importance of accounting for residual spatial correlation.
@@
Figure 5: Pre-trends test using 2017 as the “treated” period; using Spatial Weighted Bootstrap with the same Matérn parameters as our main analysis.
@@
As shown in Figure 5, we did not observe an effect when using 2017 as the “treated” period (to check for pre-trends / policy anticipation). Using 2016 & 2017 as the “treated” period (Figure S8) similarly shows no evidence of anticipation.
Lastly, including GPW site area as a covariate in our sensitivity analysis (Figure S9) did not substantially change the results.
4.2 Complementary Exploratory Data Analysis Results
@@ 5 Discussion
In this case study, we observed a 3.4% increase in PM2.5 at dump sites in Indonesia in 2018-2019 compared to 2012-2017. This effect is on a similar scale to the results of Unfried and Wang (2022), who assessed the domestic impacts of China’s plastic waste import ban and estimated an average reduction of 3.7 $\mu g/m^{3}$, corresponding to 8.5% reduction of overall PM2.5 concentrations in China as a result of their 2018 ban. It makes sense that our observed effect would be smaller, given that Indonesia was only one of the countries to which waste that would otherwise have gone to China was diverted. Unfried and Wang (2022) similarly did not observe an anticipation effect in mid-2017, when China announced its upcoming ban to the World Trade Organization.
@@
Beyond the setting of plastic waste and air quality, this paper offers two important tools that span a wide range of quasi-experimental studies: (1) a robust method to estimate the impact of a “universal” intervention (such that there are no control locations) as a function of a continuous exposure and (2) a generalizable inferential procedure for spatially correlated data in the context of a continuous exposure.
In the interest of helping practitioners understand these tools, we note that interpretation / inference for an effect curve like ours is inherently more complicated than for the effect of a binary treatment: we are interested in both the magnitude and the shape of the curve, to assess “compatibility” (Rafi and Greenland, 2020) with our hypothesis that air pollution in Indonesia increased in the wake of China’s 2018 ban, heterogeneously with respect to port proximity. A more nuanced aspect is that the plausibility of the curve shape influences our interpretation of the magnitude. In this analysis, our findings of no effect at low values of port proximity, a substantial effect at medium-high values of port proximity, and a weaker/suggestive effect at very-high values of port proximity are very plausible and compatible with our domain knowledge. To be clear, dose values whose pointwise confidence interval for the ADT excludes zero (the null) may be interpreted as doses for which there is some evidence of a treatment effect; however, the use of uniform confidence bands may be preferred for formal inferential purposes to avoid multiple comparisons concerns or to test hypotheses regarding the functional form of the effect curve. Hettinger et al. 2025a note in their Web Appendix F that uniform confidence intervals could potentially be obtained using a combination of cross-fitting and scaling of the bootstrapped standard errors by a constant specific to both the kernel and the bandwidth parameter used in LLKR and the range of the dose variable, which is an extension of Lee et al. (2017). However, as mentioned in our Section 3.5, cross-fitting in a spatiotemporal setting like ours is an active area of research.
@@
Second, while our interest is primarily on air pollution generated by the burning of waste, it is possible that some of the observed increase in PM2.5 at open dump sites is attributable to the transportation of this waste, for instance from diesel trucks making more trips to dump sites and, for sites close to the coast, from cargo ships. Our complementary (exploratory) analysis illustrating the uptick in active fires overlapping with dump sites offers some evidence that at least part of the observed effect is due to increases in waste burning. Unfortunately, speciated PM2.5 estimates, which might enable the disentangling of emissions from waste burning and transportation, tend to be available only at coarse spatiotemporal resolutions globally and rely more heavily on models (European Commission, Joint Research Centre, 2023), which is likely to frustrate the application of quasi-experimental statistical methods. For smaller scale studies, Islam et al. (2022) demonstrated the viability of measuring a molecular tracer of plastic burning with in-situ monitors. Alternatively, Scott et al. (2023) applied machine learning-based image classification to identify smoke plumes in high-resolution (3m x 3m) satellite imagery over the largest municipal landfill in the Maldives. However, in addition to being computationally intensive, this approach is not directly translatable into air pollution concentrations, which are easier to link to the rich literature on health impacts, and hence tend to be more relevant for policy makers.
@@
A more general limitation of our approach is that we cannot estimate the effects of distributed waste burning, i.e., at the household level or at communal dump sites that were too small to be detected by GPW. Future research, and likely investment in ground-level air quality monitors (and/or lower-cost sensors), will be needed to estimate these distributed burning emissions, which may ultimately pose greater risks to public health due to higher frequency of burning, greater proximity to people, and lower dispersive dilution of air pollution at the ground level (Pathak et al., 2024) compared to taller plumes of smoke resulting from larger fires at larger dump sites.
Despite its limitations, this study provides evidence that the export of plastic waste to Indonesia, and subsequent open burning of import-related waste in Indonesia, resulted in higher concentrations of PM2.5 – at least near large open dump sites. This adds to the bodies of work documenting (a) air quality impacts of plastic waste and (b) environmental degradation tied to the global waste trade. Policy interventions to address open burning of plastic waste will have to contend with the phenomenon that simply prohibiting open burning is seldom effective (Pathak et al., 2024), though burning at major sites may be easier to monitor and restrict (Scott et al., 2023). In terms of the plastic waste trade: in 2021, Indonesia restricted the import of non-hazardous waste to 15 specific ports (Septiono and Ismawati, 2022) and, in 2025, banned their import of plastic waste altogether (Neilson, 2025). In mid-2025, Malaysia followed suit, only allowing plastic waste from parties that have ratified the Basel Convention on the Control of Transboundary Movements of Hazardous Wastes and their Disposal (SIRIM QAS International, 2025). The effectiveness of these and future policies, at least for improving air quality but potentially also for other kinds of environmental outcomes, can in turn be evaluated using methods similar to those presented in this paper.
6 Competing interests
No competing interest is declared.
7 Author contributions statement
E.M.C. conceptualized this project and compiled the dataset. E.M.C. and R.C.N. developed the statistical method. E.M.C. wrote the code and carried out the analysis. E.M.C. drafted the manuscript, and both E.M.C. and R.C.N. edited the manuscript. E.M.C. led the revisions in response to reviewer comments, with feedback from R.C.N.
8 Acknowledgements
This work was supported by an NSF Graduate Research Fellowship (EMC), NIH award K01ES032458 (RCN), and the Harvard Climate Change Solutions Fund (RCN). The computation for this paper was performed on (a) the FASRC Cannon cluster supported by the FAS Division of Science Research Computing at Harvard University and (b) the Alpine high performance computing resource at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University and with support from NSF grants OAC-2201538 and OAC-2322260.
@@ Appendix A Estimation Details
A.1 Generalized Propensity Score (GPS) Estimation
@@
After fitting the initial model for $D$, we fit a model to the residuals $D_{i}-\hat{D}_{i}$ given $\mathbf{X}$, within $B=1$. (For efficiency, we use the same hyperparameters for this residual model as for the initial model of $D$.) Then, we apply a kernel density estimator to $(D_{i}-\hat{D}_{i})/\hat{\epsilon}_{i}$ to obtain $\pi_{D}$. For $f(D|B=1)=\int_{\mathbb{X}}\pi_{D}(D|B=1,\mathbf{x})dP(\mathbf{x}|B=1)$, we interpolate over the observed $\mathbf{X}$ for specified values of $D$.
@@ Appendix B Supplementary Figures and Tables
Figure S5: Directed Acyclic Graph (DAG) illustrating our application. Unobserved variables are shown in gray, while observed variables are shown in black. Interactions are suggested by intersecting arrows, as proposed by Weinberg (2007). Note that this is a slightly simplified DAG, intended to motivate our use of port proximity, rather than all the remotely sensed proxies we utilize.
Figure S6: The confounder-naive results; using Spatial Weighted Bootstrap with the same Matérn parameters as our main analysis.
Figure S7: Comparing the results using Non-Spatial and Spatial Weighted Bootstrap. Note that the point estimates are the same.
Figure S8: Pre-trends test for 2016-2017; using Spatial Weighted Bootstrap with the same Matérn parameters as our main analysis.
Figure S9: Sensitivity analysis including GPW site area as a covariate; using Spatial Weighted Bootstrap with the same Matérn parameters as our main analysis.
@@ Appendix C Simulations Exploring the Spatial Weighted Bootstrap
In this appendix, our primary aim is to evaluate the characteristics of the confidence intervals generated by our Spatial Weighted Bootstrap procedure, relative to those generated by a Non-spatial Weighted Bootstrap. Here, calculation of the ADT curve’s bias and mean absolute error (MAE) is an intermediate step to better understand the confidence intervals, as Hettinger et al. (2025a) present extensive investigation of a very similar ADT estimator.
@@ C.1 Simulating the Outcome
$\tilde{Y}_{it}=\mu_{it}+\sigma_{\text{sp}}\gamma_{i}+\sigma_{\text{ind}}\epsilon_{it}$
where $\gamma_{i}\sim MVN(0,\Sigma)$, where the correlation function to generate the entries of $\Sigma$ is Corr(m;L)=exp(m/L)Corr(m;L)=exp(-m/L) [Exponential form] and $\epsilon_{it}\sim N(0,1)$. The parameters $L$ (the true spatial range), $\sigma_{\text{sp}}$, and $\sigma_{\text{ind}}$ are constants, which we specify to obtain four distinct simulation scenarios:
@@
Note that both Scenario 2 (increasing $L$) and Scenario 3 (increasing $\sigma_{\text{sp}}$) generate greater similarity in the value of outcomes that are spatially proximate, through different mechanisms, relative to Scenario 1. Whereas, Scenario 4 simply increases the random noise in the outcomes.
@@ C.2 Estimating the Effect Curve and Confidence Interval
When implementing our ADT estimation procedure on each simulated dataset, we correctly specify the outcome models’ mean function (just the variables, not the coefficients) and estimate the propensity score models $\pi_{B}$ and $\pi_{D}$ using Ranger (the same algorithm as in our main analysis). Cross-fitting is again used for $\mu_{1}$ and $\pi_{D}$.
When applying the Spatial Weighted Bootstrap, we fit both Exponential and Power covariance functions (with corresponding correlation functions Corr(m;L)=exp(m/L)Corr(m;L)=exp(-m/L) and $Corr(m;a,z)=1-zm^{a}$, respectively) to the empirical variogram of the residuals from the LLKR stage and use whichever model has the smaller RMSE to generate the spatial weights. As noted in Section 3.4.2, the spatial structure of the LLKR residuals is different than the spatial structure of the outcome (here, $\tilde{Y}$), so traditional notions of correctly vs incorrectly specifying the spatial correlation model do not directly apply here.